OpenStudy (anonymous):

Need help on this problem please Right now, Simran's father is three times as old as Simran. Five years ago, he was four times as old as Simran. How old is each person right now?

OpenStudy (***[isuru]***):

hmmm....... In the precent if the age of simran = "x" then his father = "3x" ( fathers' age is three time as simrans ) Five years ago.. which mean simran = "x -5" father = "3x -5" since father was 4 time old as simran 3x - 5 = 4 ( x -5) now if u solve this for "x" ... u will get the age of simran in the present.. multiply it by 3.. u will get the age of his father.. got it :) ?